OpenStudy (anonymous):

If the length of each side of a cube is doubled, is the surface area doubled? Use an example to support your answer

OpenStudy (anonymous):

like my new pic oh and help

OpenStudy (anonymous):

yeah it nice

OpenStudy (anonymous):

thanks lol

OpenStudy (anonymous):

answer:no that can't be

OpenStudy (anonymous):

For example "2"x2"x2" cube -- will have a surface area on each side of 4"sq (2x2=4) --- now make that 4"x4"x4" -- this cube will have a surface area on each side of 16"sq (4x4=16)"

OpenStudy (anonymous):

one mor thing r u in the pic and thanks

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

who r u

OpenStudy (anonymous):

on the right

OpenStudy (anonymous):

u r pretty well got to go thanks for everything

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you're most welcome sis ttyl :)

OpenStudy (anonymous):

dpalnc

OpenStudy (anonymous):

lets see if a cube has edge length of x then the surface area is A=6x^2. if the edges are doubled, then the surface area is 6*(2x)^2 = 6*2x*2x = 4*6x^2 = 4A it looks like it is 4 times bigger.... :)